OpenStudy (anonymous):

solve the equation by competing the square. if necessary, round the nearest hundredth. x^2 - 18x = 19 a. 1; 19 b. -1; 19 c. 3; 6 d. -3; 1

OpenStudy (anonymous):

@amorfide

OpenStudy (amorfide):

okay so to complete the square \[Ax ^{2} + Bx = C\] make it equal to 0 \[Ax ^{2} + Bx + C = 0\] completing the square has a formula \[(x+\frac{ B }{ 2 }) - (\frac{ B }{ 2 })^{2} + C = 0\]

OpenStudy (amorfide):

oops

OpenStudy (amorfide):

\[(x+\frac{ B }{ 2 })^{2} - (\frac{ B}{ 2 })^{2} + C =0\]

OpenStudy (amorfide):

sorry I forgot to square it in my first post

OpenStudy (amorfide):

can you do this?